OpenStudy (anonymous):

((-5+sqrt13)/2)y+(-5-sqrt13)/2=0 I need to solve for Y.

OpenStudy (anonymous):

please anyone I dont know how to get started

OpenStudy (anonymous):

do you know how to solve this

OpenStudy (anonymous):

please

OpenStudy (mertsj):

\[\frac{-5+\sqrt{13}}{2}y+\frac{-5-\sqrt{13}}{2}=0\]

OpenStudy (mertsj):

Do I have it right?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Subtract the second fraction from both sides.

OpenStudy (anonymous):

Yes that is when I get the problem. That step and the one that follows

OpenStudy (mertsj):

\[\frac{-5+\sqrt{13}}{2}y=- \frac{-5-\sqrt{13}}{2}=\frac{5+\sqrt{13}}{2}\]

OpenStudy (mertsj):

Now multiply both sides of the equation by the reciprocal of the coefficient of y

OpenStudy (anonymous):

Ok, when i do that for some reason everything cancels out

OpenStudy (mertsj):

\[y=\frac{5+\sqrt{13}}{2} \times \frac{2}{-5+\sqrt{13}}=\frac{\sqrt{13}+5}{\sqrt{13}-5}\]

OpenStudy (anonymous):

ok Im following

OpenStudy (mertsj):

Now rationalize the denominator by multiplying numerator and denominator by sqrt13+5

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

\[y=\frac{\sqrt{13}+5}{\sqrt{13}-5}\times\frac{\sqrt{13}+5}{\sqrt{13}+5}=\frac{13+10\sqrt{13}+25}{13-25}=\frac{38+10\sqrt{13}}{-12}\]

OpenStudy (mertsj):

Now divide out the common factor 2 from numerator and denominator.

OpenStudy (anonymous):

ok following along

OpenStudy (mertsj):

Can you divide out the 2?

OpenStudy (anonymous):

no :( Im sorry I have not done this in a long time.

OpenStudy (mertsj):

What is 38 divided by 2?

OpenStudy (anonymous):

19

OpenStudy (mertsj):

ok. Replace the 38 with 19. What is 10sqrt13 divided by 2?

OpenStudy (anonymous):

5sqrt13?

OpenStudy (mertsj):

yes. Replace the 10sqrt13 with 5sqrt13. What is -12 divided by 2?

OpenStudy (anonymous):

-6

OpenStudy (mertsj):

Yes. replace the -12 with -6 and now you are done.

OpenStudy (anonymous):

oh ok, for some reason the answer I have is (5sqrt 13 +13)/26 ?

OpenStudy (anonymous):

It makes since when I followed along with you, but the answer does not make since?

OpenStudy (mertsj):

Would you please check and make sure the problem is posted exactly as it is stated in your book?

OpenStudy (anonymous):

yes, it could be wrong in the book

OpenStudy (anonymous):

You have helped me alot thank you so much:)

OpenStudy (mertsj):

yw